OpenStudy (anonymous):

lim x tends to 0 (cosx+sinx)whole power 1/x

hartnn (hartnn):

@harishk8 Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) you can let L = \(\lim \limits{x\rightarrow 0 (\cos x+\sin x)^{1/x}}\) then take, log on both sides, what you get ? log L =.... ?